# Exploring Graph-Transformer Out-of-Distribution Generalization Abilities

### v1

@@ 4.3 Vanilla GT/GPS versus MPNNs equipped with DG algorithms (RQ1)
Figure 2 clearly shows that GPS (green), achieves the highest OOD accuracy across all benchmarks, even when trained with standard ERM (without explicit DG algorithms). In contrast, vGIN (blue), even when equipped with DG methods like CORAL and DANN, consistently underperforms compared to GPS. Particularly on the challenging CMNIST-color task, GPS stands out, whereas MHA (red) slightly surpasses it on Motif tasks but does not match its overall robustness across all benchmarks. For example, GPS outperforms vGIN by a large margin on CMNIST-color (87% vs. 29.8%) and remains highly competitive on all other benchmarks, with an average OOD accuracy of 86.6% compared to 55.6% for vGIN and 70.6% for MHA.
These results underscore the benefits of using GT for improving OOD robustness (as also observed in computer vision [18, 20]). The ability of global attention and context modeling to capture long-range dependencies appears to significantly boost OOD accuracy under distribution shifts on those tasks. Notably, hybrid architectures such as GPS, which integrate local message passing with global attention, further amplify this advantage. These findings indicate that incorporating GT, particularly in a hybrid backbone, represents a promising direction for advancing OOD generalization in graph learning.
@@ 4.4 GT/GPS versus MPNNs when equipped with DG algorithms (RQ2)
Table 1 benchmarks three backbone architectures (vGIN, MHA, and GPS) across four distribution-shifted datasets from the GOOD benchmark under seven DG algorithms. While vGIN often performs well in-domain, it struggles under distribution shifts.
@@
In the Motif with basis shift task, MHA and GPS exhibit closely matched OOD accuracy and gap, suggesting that both attention-based and hybrid backbones are well suited to handling structural shifts in graph topology.
@@
Last, for the SST2 with length shift task, GPS achieves the highest OOD accuracy (81.05%) with CORAL, while also maintaining low accuracy gaps across DG algorithms. Although vGIN attains relatively strong OOD performance here compared to other tasks, the gains from GPS remain consistent and underline its robustness.
@@
Across various graph classification tasks and DG algorithms, GPS consistently achieves the highest or near-highest OOD accuracies and smallest accuracy GAP, demonstrating the advantage of combining local message passing with global attention. In contrast, vGIN exhibits large performance drops, especially on CMNIST and Motif—where the accuracy gaps exceed 35 to 50 points, highlighting its sensitivity to distribution shifts. Meanwhile, MHA perform well on Motif tasks but less effectively on others. These patterns suggest that GPS offers a more reliable backbone for maintaining performance under distribution shift. The results are visualized in Figure 3, where the left panel shows the OOD test accuracy results, and the right one presents the gap accuracies. ID results are discussed in Appendix B.
@@
DG Alg. Backbone CMNIST-color Motif-size Motif-basis SST2-length
ID acc. OOD acc. Gap ID acc. OOD acc. Gap ID acc. OOD acc. Gap ID acc. OOD acc. Gap
ERM vGIN 78.84 28.74 50.11 90.93 37.23 53.70 92.94 47.20 45.73 89.82 69.57 20.25
MHA 46.38 23.43 22.95 92.97 88.34 4.63 93.62 92.48 1.13 86.20 78.23 7.97
GPS 89.32 87.03 2.29 92.96 84.24 8.72 93.54 92.17 1.38 86.55 80.86 5.70
Mixup vGIN 78.79 26.05 52.75 90.02 39.63 50.39 92.91 48.99 43.92 89.74 70.16 19.58
MHA 52.88 26.58 26.30 92.74 81.73 11.01 93.58 92.03 1.56 86.47 77.86 8.61
GPS 92.81 56.41 36.41 92.88 74.31 18.58 93.57 86.17 7.40 86.64 79.20 7.44
DANN vGIN 76.24 29.85 46.39 92.83 33.93 58.89 92.98 58.72 34.25 89.93 77.64 12.29
MHA 50.91 22.68 28.23 92.90 84.94 7.96 93.59 92.56 1.03 86.47 79.52 6.96
GPS 89.95 82.69 7.26 92.81 74.02 18.79 93.60 90.29 3.31 86.67 80.40 6.27
Coral vGIN 78.94 26.09 52.85 93.07 54.20 38.88 92.90 54.47 38.43 89.83 74.26 15.57
MHA 53.57 41.10 12.47 92.99 84.46 8.53 93.60 89.41 4.19 86.66 76.15 10.51
GPS 90.53 84.03 6.49 92.84 78.33 14.52 93.52 92.18 1.34 87.21 81.05 6.16
GroupDRO vGIN 78.06 28.71 49.35 92.89 39.91 52.98 92.97 43.65 49.32 89.66 73.11 16.54
MHA 48.25 37.10 11.15 92.96 89.67 3.29 93.58 92.78 0.80 86.33 77.99 8.35
GPS 89.70 82.96 6.74 92.87 81.39 11.48 93.61 89.14 4.47 86.54 80.69 5.85
IRM vGIN 77.99 28.13 49.86 92.45 35.93 56.53 92.95 40.97 51.99 89.75 76.76 12.98
MHA 50.74 26.61 24.13 92.99 85.82 7.18 93.59 92.34 1.25 86.48 77.89 8.58
GPS 89.95 89.26 0.68 92.93 82.93 10.00 93.60 92.24 1.36 86.69 80.03 6.66
VREx vGIN 78.54 26.04 52.50 91.30 37.01 54.29 85.83 35.10 50.73 87.54 64.96 22.59
MHA 46.28 28.76 17.52 92.71 87.09 5.63 77.87 64.95 12.92 85.43 75.48 9.95
GPS 88.28 87.72 0.56 88.87 65.14 23.73 80.47 67.97 12.50 84.47 78.75 5.73
@@ 4.5 Quantification of model generalization abilities (RQ3)
Table 2 shows notable differences across backbones and DG algorithms, with backbone choice often exerting a stronger influence than the algorithm itself. Interestingly, ERM paired with MHA or GPS frequently matches or outperforms specialized DG approaches. At the same time, performance is highly dataset-dependent. On CMNIST-color, GPS achieves the best Silhouette score (0.216 with IRM) and lowest MMD distance (0.001, tied across multiple algorithms), while MHA and vGIN produce negative Silhouette scores for all DG methods, indicating poor domain separation. For both Motif tasks, MHA ranks highest (Silhouette score $>0.6$ even with ERM; MMD distance $<0.02$), while GPS also performs well across algorithms, and vGIN struggles consistently. For SST2-length, GPS with Mixup achieves the best Silhouette score (0.372), and MHA with GroupDRO reaches the lowest MMD distance (0.009), while vGIN exhibits weak separation (e.g., 0.071 Silhouette score with VREx). These trends underscore that the architectural design of MHA and GPS leads to more robust, tighter clusters and domain-aligned representations, often surpassing what DG-specific methods can achieve with weaker backbones. Additionally, we observe that optimizing for Silhouette score and MMD distance does not always coincide, highlighting the importance of evaluating these two complementary aspects, clustering and alignment, when assessing domain generalization.
@@
DG Alg. Backbone CMNIST-color Motif-size Motif-basis SST2-length
Silhouette MMD Silhouette MMD Silhouette MMD Silhouette MMD
ERM vGIN -0.079 0.060 -0.011 0.353 0.149 0.235 0.223 0.156
MHA -0.053 0.203 0.682 0.008 0.608 0.027 0.337 0.018
GPS 0.197 0.005 0.483 0.046 0.577 0.052 0.328 0.025
Mixup vGIN 0.048 0.123 -0.039 0.287 0.047 0.366 0.319 0.125
MHA -0.013 0.135 0.566 0.051 0.470 0.036 0.341 0.015
GPS 0.198 0.073 0.382 0.057 0.368 0.076 0.372 0.058
DANN vGIN -0.129 0.047 -0.026 0.352 -0.026 0.159 0.143 0.121
MHA -0.109 0.038 0.551 0.015 0.379 0.053 0.289 0.017
GPS -0.070 0.001 0.255 0.042 0.403 0.068 0.237 0.034
Coral vGIN -0.015 0.087 0.056 0.166 0.272 0.238 0.223 0.116
MHA -0.003 0.077 0.630 0.016 0.614 0.052 0.321 0.018
GPS 0.195 0.025 0.376 0.052 0.452 0.094 0.345 0.055
GroupDRO vGIN -0.063 0.069 -0.035 0.238 -0.010 0.297 0.165 0.120
MHA -0.022 0.232 0.666 0.011 0.615 0.017 0.327 0.009
GPS 0.195 0.017 0.400 0.040 0.537 0.066 0.337 0.034
IRM vGIN -0.045 0.075 -0.016 0.217 0.112 0.347 0.155 0.213
MHA -0.040 0.177 0.640 0.013 0.570 0.033 0.300 0.013
GPS 0.216 0.001 0.396 0.042 0.565 0.040 0.324 0.021
VREx vGIN -0.070 0.078 0.066 0.240 -0.018 0.428 0.071 0.268
MHA -0.043 0.098 0.551 0.018 0.078 0.078 0.187 0.125
GPS 0.198 0.001 0.171 0.064 0.151 0.224 0.043 0.164
@@ 4.6 Complementary abilities of model evaluation and generalization metrics (RQ4)
Although OOD accuracy, ID-OOD accuracy gap, MMD distance, and Silhouette score often trend in the same direction, they probe different aspects of generalization and therefore do not always correlate tightly. Models that attain similar OOD accuracy can differ markedly in the way they organize their latent representations. For example, in the Motif-basis experiments, both MHA-VREx and GPS-VREx exhibit a $gap=12$, yet their OOD accuracies diverge by more than three points. A similar pattern occurs on SST2: vGIN’s DANN and IRM variants yield only slightly lower OOD accuracy than MHA and GPS, but their accuracy gap is substantially larger.
Moreover, choosing a model solely on OOD accuracy and accuracy gap can also be misleading. For example, in the case of Motif-basis, the OOD accuracy and accuracy gap of MHA-ERM and GPS-ERM appear close to those of MHA-DANN and GPS-CORAL, respectively; however, MHA-DANN produces a markedly lower Silhouette score—signaling poorly separated class clusters—while GPS-CORAL shows a much higher MMD distance, indicating misaligned domain feature distributions.
In most cases, as intuition suggests, high OOD accuracy is typically accompanied by a small accuracy gap. It is also associated with a low MMD distance, which indicates well-aligned domains, and a high Silhouette score, which reflects well-separated class clusters. Yet the counterexamples above illustrate that these regularities are not universal. Reporting all four metrics offers a more comprehensive and interpretable understanding of each model’s generalization capabilities. As a result, better-informed decisions can be made when selecting models for real-world applications.
@@ 5 Conclusion
This work presents the first systematic evaluation of Graph Transformer (GT) backbones for out-of-distribution (OOD) generalization in supervised graph classification. Our results demonstrate that GraphGPS (GPS), a hybrid architecture integrating local message passing with global transformer attention, consistently outperforms both MPNN (vGIN) and pure attention-based (MHA) models, underscoring the importance of backbone design for OOD robustness. Notably, GPS achieves an average OOD accuracy of 86.6% across benchmarks—substantially higher than MHA (70.6%) and vGIN (55.6%), even when the latter is equipped with domain generalization algorithms. On challenging shifts such as CMNIST-color, GPS attains 87% OOD accuracy compared to vGIN’s 29.8%.

### v2

@@ 4.3 Vanilla GT/GPS versus MPNNs equipped with DG algorithms (RQ1)
Figure 2 clearly shows that GPS (green) achieves the highest average OOD accuracy across benchmarks, even when trained only with standard ERM (without explicit DG algorithms). In contrast, vGIN (blue), although equipped with DG algorithms, underperforms compared to both GPS and MHA on CMNIST, Motif, and SST tasks. While vGIN leads on both HIV’s tasks, it can’t be seen in Figure 2 but can be seen in Table 1, this gap narrows down in the HIV-size but not in HIV-scaffold. vGIN leads on both HIV tasks, however, since Figure 2 reflects vGIN equipped with DG algorithms while GPS and MHA are vanilla, the actual gap narrows on HIV-size but remains pronounced on HIV-scaffold, as shown in Table 1. Particularly on the challenging CMNIST-color task, GPS stands out, whereas MHA (red) slightly surpasses it on Motif tasks but does not match its overall robustness across all benchmarks. For example, GPS outperforms vGIN by a large margin on CMNIST-color (87% vs. 29.8%) and remains highly competitive on all other benchmarks, with an average OOD accuracy of 67.3% compared to 52% for vGIN and 56.7% for MHA.
These results underscore the benefits of using GT for improving OOD robustness (as also observed in computer vision [19, 21]). The ability of global attention and context modeling to capture long-range dependencies appears to significantly boost OOD accuracy under distribution shifts on those tasks. Notably, hybrid architectures such as GPS, which integrate local message passing with global attention, further amplify this advantage. At the same time, results on specific tasks such as HIV-scaffold indicate that in certain shifts, local message passing alone may be better suited. These findings indicate that incorporating GT, particularly in a hybrid backbone, represents a promising direction for advancing OOD generalization in graph learning.
@@ 4.4 GT/GPS versus MPNNs when equipped with DG algorithms (RQ2)
Table 1 benchmarks three backbone architectures (vGIN, MHA, and GPS) across six distribution-shifted tasks from the GOOD benchmark under seven DG algorithms. While vGIN often performs well in-domain, it struggles under distribution shifts, at four out of six tasks.
@@
In the Motif with basis shift task, MHA and GPS exhibit closely matched OOD accuracy and gap, suggesting that both attention-based and hybrid backbones are well-suited to handling structural shifts in graph topology.
@@
For the SST2 with length shift task, GPS achieves the highest OOD accuracy (83.29%) with GroupDRO, while also maintaining low accuracy gaps across DG algorithms. Although vGIN attains relatively strong OOD performance here, the gains from GPS remain consistent and underline its robustness.
Last for the HIV tasks, when comparing the strongest OOD variants of each backbone, vGIN leads by roughly 2% on the size shift and 5% on the scaffold shift, suggesting that locality in message passing can be better suited for certain distribution shifts.
@@
Across four out of six graph classification tasks and various DG algorithms, GPS consistently achieves the highest or near-highest OOD accuracies and smallest accuracy GAP, demonstrating the advantage of combining local message passing with global attention. In contrast, vGIN exhibits large performance drops (except for the HIV tasks), especially on CMNIST and Motif—where the accuracy gaps exceed 25 to 50 points, highlighting its sensitivity to distribution shifts. Meanwhile, MHA perform well on Motif tasks but less effectively on others. These patterns suggest that GPS offers a more reliable backbone for maintaining performance under most distribution shifts. The results are visualized in Figure 3, where the left panel shows the OOD test accuracy results, and the right one presents the gap accuracies. ID results are discussed in Appendix B.
@@
DG Alg. Backbone CMNIST-color Motif-size Motif-basis SST2-length HIV-size HIV-scaffold
ID acc. OOD acc. Gap ID acc. OOD acc. Gap ID acc. OOD acc. Gap ID acc. OOD acc. Gap ID acc. OOD acc. Gap ID acc. OOD acc. Gap
ERM vGIN 78.84 28.74 50.11 92.15 49.21 42.94 92.63 66.59 26.05 89.65 78.61 11.04 82.65 61.34 21.30 80.77 69.51 11.26
ERM MHA 46.38 23.43 22.95 92.97 88.34 4.63 93.62 92.48 1.13 86.20 78.24 7.97 71.86 53.94 17.92 70.24 60.61 9.63
ERM GPS 89.32 87.03 2.29 92.96 84.24 8.72 93.54 92.17 1.38 88.71 82.80 5.92 81.55 58.05 23.51 80.56 67.05 13.51
Mixup vGIN 78.80 26.05 52.75 91.80 48.60 43.20 92.62 64.56 28.06 90.02 79.66 10.36 82.37 62.95 19.42 80.72 69.22 11.50
Mixup MHA 52.88 26.58 26.30 92.74 81.73 11.01 93.58 92.03 1.56 86.47 77.86 8.61 68.83 56.67 12.16 66.08 57.69 8.39
Mixup GPS 92.81 56.41 36.41 92.88 74.31 18.58 93.57 86.17 7.40 89.06 82.14 6.91 80.30 53.67 26.63 77.10 61.77 15.33
DANN vGIN 76.24 29.85 46.39 91.99 46.39 45.60 92.58 63.33 29.25 89.76 80.20 9.56 82.34 62.08 20.25 70.93 60.34 10.60
DANN MHA 50.91 22.68 28.23 92.90 84.94 7.96 93.59 92.56 1.03 86.47 79.52 6.96 69.08 55.21 13.87 60.80 56.36 4.45
DANN GPS 89.95 82.69 7.26 92.81 74.02 18.79 93.60 90.29 3.31 89.14 82.87 6.27 80.99 57.15 23.84 64.54 59.28 5.26
Coral vGIN 78.94 26.09 52.85 92.07 49.83 42.24 92.64 66.80 25.84 89.49 78.05 11.44 83.08 61.41 21.66 82.12 69.72 12.40
Coral MHA 53.57 41.10 12.47 92.99 84.46 8.53 93.60 89.41 4.19 86.66 76.15 10.51 69.38 55.43 13.95 68.81 59.25 9.56
Coral GPS 90.53 84.03 6.49 92.84 78.33 14.52 93.52 92.18 1.34 88.90 82.93 5.97 81.35 55.64 25.71 80.47 66.79 13.68
GroupDRO vGIN 78.06 28.71 49.35 91.99 52.24 39.76 92.60 66.92 25.68 89.53 77.83 11.70 82.64 61.67 20.97 80.61 68.92 11.69
GroupDRO MHA 48.25 37.10 11.15 92.96 89.67 3.29 93.58 92.78 0.80 86.33 77.99 8.35 68.96 57.44 11.53 66.84 60.52 6.32
GroupDRO GPS 89.70 82.96 6.74 92.87 81.39 11.48 93.61 89.14 4.47 88.72 83.29 5.43 81.37 59.04 22.33 77.06 62.56 14.50
IRM vGIN 77.99 28.13 49.86 91.97 52.20 39.77 92.62 66.62 26.01 89.47 77.68 11.80 72.81 53.23 19.58 80.90 72.06 8.84
IRM MHA 50.74 26.61 24.13 92.99 85.82 7.18 93.59 92.34 1.25 86.48 77.89 8.59 62.26 58.14 4.12 68.38 58.57 9.81
IRM GPS 89.95 89.26 0.68 92.93 82.93 10.00 93.60 92.24 1.36 88.89 82.23 6.66 70.59 60.58 10.02 74.01 62.89 11.12
VREx vGIN 78.54 26.04 52.50 87.82 41.79 46.03 89.65 45.38 44.28 89.17 80.32 8.85 81.31 61.15 20.16 75.47 66.51 8.96
VREx MHA 46.28 28.76 17.52 92.71 87.09 5.63 77.87 64.95 12.92 85.43 75.48 9.95 66.20 56.03 10.17 62.50 57.66 4.84
VREx GPS 88.28 87.72 0.56 88.87 65.14 23.73 80.47 67.97 12.50 86.25 78.58 7.67 77.92 57.46 20.46 69.10 59.23 9.87
@@ 4.5 Quantification of model generalization abilities (RQ3)
Table 2 shows notable differences across backbones and DG algorithms, with backbone choice often exerting a stronger influence than the algorithm itself. Interestingly, ERM paired with MHA or GPS frequently matches or outperforms specialized DG approaches. At the same time, performance is highly dataset-dependent. On CMNIST-color, GPS achieves the best Silhouette score (0.216 with IRM) and lowest MMD distance (0.001, tied across multiple algorithms), while MHA and vGIN produce negative Silhouette scores for all DG methods, indicating poor domain separation. For both Motif tasks, MHA ranks highest (Silhouette score >0.68>0.68 even with ERM; MMD distance <0.08<0.08), while GPS also performs well across algorithms, and vGIN struggles consistently. For SST2-length, GPS with Mixup achieves the best Silhouette score (0.356), and MHA with Mixup reaches the lowest MMD distance (0.012). For both HIV tasks, vGIN have the highest Silhouette scores, but MHA have the lowest MMD distances. These trends (on four out of six benchmarks) underscore that the architectural design of MHA and GPS leads to more robust, tighter clusters and domain-aligned representations, often surpassing what DG-specific methods can achieve with weaker backbones. Additionally, we observe that optimizing for Silhouette score and MMD distance does not always coincide, highlighting the importance of evaluating these two complementary aspects, clustering and alignment, when assessing domain generalization.
@@
DG Alg. Backbone CMNIST-color Motif-size Motif-basis SST2-length HIV-size HIV-scaffold
Silh. MMD Silh. MMD Silh. MMD Silh. MMD Silh. MMD Silh. MMD
ERM vGIN -0.060 0.071 0.038 0.197 0.111 0.310 0.240 0.046 0.092 0.124 0.145 0.020
MHA -0.062 0.312 0.685 0.007 0.758 0.008 0.309 0.016 -0.117 0.062 -0.029 0.022
GPS 0.094 0.136 0.429 0.053 0.602 0.060 0.213 0.077 0.087 0.035 0.128 0.010
Mixup vGIN 0.058 0.084 0.031 0.189 0.130 0.263 0.348 0.086 0.284 0.094 0.307 0.025
MHA -0.080 0.266 0.545 0.049 0.712 0.021 0.354 0.012 -0.020 0.025 -0.145 0.015
GPS 0.194 0.099 0.371 0.061 0.596 0.048 0.356 0.050 0.135 0.036 0.156 0.015
DANN vGIN -0.173 0.050 -0.003 0.236 -0.008 0.553 0.111 0.051 0.312 0.023 0.281 0.127
MHA -0.174 0.214 0.530 0.023 0.502 0.066 0.290 0.022 -0.050 0.005 0.051 0.008
GPS -0.056 0.064 0.207 0.057 0.455 0.062 0.148 0.108 0.313 0.032 0.013 0.005
Coral vGIN -0.017 0.105 0.041 0.205 0.104 0.322 0.276 0.070 0.233 0.078 0.258 0.022
MHA -0.011 0.097 0.616 0.011 0.714 0.024 0.299 0.013 -0.019 0.026 0.018 0.041
GPS 0.193 0.028 0.386 0.044 0.665 0.040 0.257 0.059 0.113 0.054 0.176 0.011
GroupDRO vGIN -0.043 0.088 0.041 0.189 0.113 0.307 0.246 0.046 0.142 0.074 0.272 0.007
MHA -0.006 0.110 0.693 0.007 0.723 0.021 0.304 0.016 -0.038 0.015 -0.115 0.010
GPS 0.190 0.026 0.368 0.047 0.633 0.066 0.324 0.046 0.115 0.035 0.128 0.009
IRM vGIN -0.032 0.077 0.039 0.202 0.112 0.309 0.237 0.046 0.336 0.284 0.150 0.013
MHA -0.045 0.305 0.635 0.013 0.730 0.026 0.267 0.016 0.088 0.017 -0.006 0.004
GPS 0.226 0.001 0.428 0.037 0.716 0.034 0.282 0.072 0.133 0.110 0.116 0.009
VREx vGIN -0.033 0.081 0.032 0.181 0.156 0.252 0.332 0.068 0.368 0.051 0.667 0.025
MHA -0.050 0.200 0.536 0.018 0.177 0.061 0.334 0.025 -0.063 0.011 0.282 0.022
GPS 0.198 0.001 0.190 0.062 0.179 0.178 0.293 0.104 0.096 0.025 0.225 0.006
@@ 4.6 Complementary abilities of model evaluation and generalization metrics (RQ4)
Although OOD accuracy, ID-OOD accuracy gap, MMD distance, and Silhouette score often trend in the same direction, they probe different aspects of generalization and therefore do not always correlate tightly. Models that attain similar OOD accuracy can differ markedly in the way they organize their latent representations. When two models achieve the same high OOD accuracy, a larger gap indicates higher ID performance, suggesting greater potential for further improving generalization, though it may also carry some risk if future OOD shifts differ from the current test set, highlighting why both OOD accuracy and the gap metric are needed to assess real-world deployment suitability.
Moreover, selecting a model solely based on OOD accuracy and accuracy gap can be misleading. For instance, in the Motif-basis task, MHA-ERM exhibits OOD accuracy and gap values similar to other variants; however, models such as MHA-DANN yield a substantially lower Silhouette score, indicating poorly separated class clusters, and a higher MMD distance, reflecting misaligned domain feature distributions, underscoring the need for complementary metrics beyond simple accuracy measures.
In most cases, as intuition suggests, high OOD accuracy is typically accompanied by a small accuracy gap. It is also associated with a low MMD distance, which indicates well-aligned domains, and a high Silhouette score, which reflects well-separated class clusters. Yet the counter examples above illustrate that these regularities are not universal. Reporting all six metrics offers a more comprehensive and interpretable understanding of each model’s generalization capabilities. As a result, better-informed decisions can be made when selecting models for real-world applications.
@@ 5 Conclusion
This work presents the first systematic evaluation of Graph Transformer (GT) backbones for out-of-distribution (OOD) generalization in supervised graph classification. Our results demonstrate that GraphGPS (GPS), a hybrid architecture integrating local message passing with global transformer attention, consistently outperforms both MPNN (vGIN) and pure attention-based (MHA) models (on four out of six benchmarks), demonstrating the importance of backbone design for OOD robustness. Notably, GPS achieves an average OOD accuracy of 67.3% across benchmarks—substantially higher than MHA (56.7%) and vGIN (52%), even when only the latter is equipped with the strongest domain generalization algorithms. On challenging shifts such as CMNIST-color, GPS attains 87% OOD accuracy compared to vGIN’s 29.8%.